(2x+ 2 x y²) dx + ( x ²) - 3y) dy = 0 solve your equation
Answer #1

(2x+2772 )dx+(x*y-3y)dy = 0 x(2+2y^)dx + y(x2 – 3) dy = 0 - 2 dx + -3 integrate dy = 0 2+2y2 1 11n | x2 – 347. in 12+33*|=0 2